# Geometry and factorization of multivariate Markov chains with applications to the swapping algorithm

### v4

@@ Abstract
This paper analyzes the factorizability and geometry of transition matrices of multivariate Markov chains. Specifically, we demonstrate that the induced chains on factors of a product space can be regarded as information projections with respect to the Kullback-Leibler divergence. This perspective yields Han-Shearer type inequalities and submodularity of the entropy rate of Markov chains, as well as applications in the context of large deviations and mixing time comparison. As a concrete algorithmic application, we introduce a projection sampler based on the swapping algorithm, which resamples the highest-temperature coordinate at stationarity at each step. We prove that such practice accelerates the mixing time by multiplicative factors related to the number of temperatures and the dimension of the underlying state space when compared with the original swapping algorithm.
Keywords: Markov chains, $f$-divergence, mutual information, Markov chain Monte Carlo, large deviations, matrix nearness problem, swapping algorithm, Han’s inequality, Shearer’s lemma, submodularity, information geometry
AMS 2020 subject classification: 60F10, 60J10, 60J22, 94A15, 94A17
@@ 1 Introduction
$\widetilde{D}_{KL}(\mu\|\nu):=\sum_{x\in\Omega}\mu(x)\ln\left(\dfrac{\mu(x)}{\nu(x)}\right),$
where the usual convention of $00\ln(0/0):=$ 0 applies. With the above setup in mind, the classical notion of mutual information between $X,Y$ is defined to be
@@
where the symbol $(\mu\otimes\nu)(x,y):=\mu(x)\nu(y)$ for all $x,y\in\Omega$ denotes the product distribution of $\mu$ and $\nu$. Note that $\mu\otimes\nu\in\mathcal{P}(\Omega^{2})$. It can be shown, via the chain rule of the KL divergence (Cover and Thomas, 2006, Theorem $2.5.3$), that $p_{X}\otimes p_{Y}$ is the unique closest product distribution to $p_{(X,Y)}$ in the sense that
$I(X;Y)=\min_{\mu,\nu\in\mathcal{P}(\Omega)}\widetilde{D}_{KL}(p_{(X,Y)}\|\mu\otimes\nu).$ (1)
In other words, the mutual information $I(X;Y)$ can broadly be interpreted as an entropic “distance” to the closest product distribution, that is, an entropic distance to independence. From (1), we immediately see that $I(X;Y)$ is non-negative, and vanishes if and only if $X,Y$ are independent (Cover and Thomas, 2006, equation $(2.90)$).
In the context of Markov chains, we consider two transition matrices $M,L$ on a common finite state space $\mathcal{X}$ and we write the set of all transition matrices on $\mathcal{X}$ to be $\mathcal{L}(\mathcal{X})$. Let $f:\mathbb{R}^{+}\to\mathbb{R}$ be a convex function with $f(1)=0$, and $\pi\in\mathcal{P}(\mathcal{X})$. Analogous to $f$-divergence between infinitesimal generators of continuous-time Markov chains in (Diaconis and Miclo, 2009, Proposition $1.5$), we define the $f$-divergence from $L$ to $M$ with respect to $\pi$ to be
@@
where $x^{i},y^{i}\in\mathcal{X}^{(i)}$ for $i=1,2$. Suppose now the state space $\mathcal{X}=\mathcal{X}^{(1)}\times\ldots\times\mathcal{X}^{(d)}$ takes on a product form for $d\in\mathbb{N}$. Given a $P\in\mathcal{L}(\mathcal{X})$, what is the closest product Markov chain? To put this question in concrete applications, we can think of an interacting particle system with $d$ particles or agents, such as the voter model (Liggett, 2004), that is described by a transition matrix $P$. What is the dynamics of the closest independent system to $P$ in which each particle or agent evolves independently of each other? That is, we are interested in seeking a minimizer of
@@
which is analogous to the classical mutual information as in (1). Note that we write $\llbracket d\rrbracket:=\{1,\ldots,d\}$ for $d\in\mathbb{N}$. We will first verify that $\mathbb{I}_{f}^{\pi}$ shares similar geometric properties with the distance to product distributions (i.e. mutual information between two random variables). As a result, $\mathbb{I}_{f}^{\pi}(P)$ can be interpreted as a distance to independence of a given $P$.
Orthogonality considerations allow us to identify and determine the closest product chain under KL divergence. This can be seen as a Markov chain version of the matrix nearness problem investigated in Li and Lim (2023); Higham (1989), as we are seeking the closest product chain from a given $P$. In the dual case when $f$ generates the reverse KL divergence, we present a large deviation principle of Markov chains where $\mathbb{I}_{f}^{\pi}(P)$ plays a role in the exponent of large deviation probability. These results are presented in Section 2.1 and 2.1.1 below.
Note that in Lacker (2023) a similar problem has been investigated in the context of diffusion processes, where the author studies the closest independent diffusion process of a given multivariate diffusion process and identifies the associated Wasserstein gradient flow and consequences for the McKean-Vlasov equation. On the other hand in the present manuscript, we focus on the closest independent Markov chain problem and the underlying geometry induced by information divergences between transition matrices.
We proceed to generalize these notions further. We introduce the leave-one-out, or more generally leave-$S$-out transition matrix, and investigate the factorizability of a transition matrix with respect to a partition or cliques of a given graph in Section 2.2 to 2.4. Observing that leave-$S$-out transition matrices are instances of Markov chain decomposition (Jerrum et al., 2004) or induced chains (Aldous and Fill, 2002), we deduce comparison results for hitting and mixing time parameters such as spectral gap and log-Sobolev constant between $P$ and its information projections.
@@
The rest of this paper is organized as follows. In Section 2, we first recall the notion of $f$-divergences between transition matrices and probability measures. We derive a few important properties of these divergences on finite product state spaces, which allow us to define an entropic distance to independence of a given multivariate $P$ in Section 2.1. In Section 2.1.1, we determine and identify the closest product chain under KL divergence, and present a large deviation principle in this context. We investigate the factorizability of $P$ with respect to partition or cliques of a given graph in Section 2.2 to 2.4. In Section 2.5, we compare mixing and hitting time parameters between $P$ and its information projections, while in Section 2.6, we show that several entropic functions that naturally arise in this paper are in fact submodular. To illustrate the applicability of projection chains, in Section 3 we propose a projection sampler and compare its mixing time with the original swapping algorithm.
@@ 2 Distance to independence and factorizability of Markov chains
On a finite state space $\mathcal{X}$, we define $\mathcal{L}=\mathcal{L}(\mathcal{X})$ as the set of transition matrices of discrete-time homogeneous Markov chains. We denote by $\mathcal{P}(\mathcal{X})$ to be the set of probability masses on $\mathcal{X}$. Let $\pi\in\mathcal{P}(\mathcal{X})$ be any given positive probability distribution, and denote $\mathcal{L}(\pi)\subset\mathcal{L}$ as the set of $\pi$-reversible transition matrices on $\mathcal{X}$, where a transition matrix $P\in\mathcal{L}$ is said to be $\pi$-reversible if $\pi(x)P(x,y)=\pi(y)P(y,x)$ for all $x,y\in\mathcal{X}$. We also say that $P\in\mathcal{L}$ is $\pi$-stationary if it satisfies $\pi P=\pi$. Suppose that $P$ is $\pi$-stationary, then the $\pi$-dual or the time reversal of $P$, $P^{*}$, is defined to be $P^{*}(x,y):=\frac{\pi(y)}{\pi(x)}P(y,x)$, for all $x,y\in\mathcal{X}$.
@@
$D_{f}^{\pi}(M\|L):=\sum_{x\in\mathcal{X}}\pi(x)\sum_{y\in\mathcal{X}}L(x,y)f\left(\dfrac{M(x,y)}{L(x,y)}\right).$ (2)
@@
$\widetilde{D}_{f}(\mu\|\nu):=\sum_{x\in\Omega}\nu(x)f\left(\frac{\mu(x)}{\nu(x)}\right),$ (3)
where we apply the usual convention that $00f(\frac{0}{0}):=$ 0 and $0f(\frac{a}{0}):=af^{\prime}(+\infty)$ with $f^{\prime}(+\infty):=\lim_{x\to 0^{+}}xf(\frac{1}{x})$ for $a>0$ in the two definitions above. We also adapt the convention that $0\cdot\infty:=0$.
In the special case of taking $f(t)=t\ln t$, we recover the KL divergence. In this case we shall write $D_{KL}^{\pi}$ and $\widetilde{D}_{KL}$ respectively. In particular, when $M,L$ are assumed to be $\pi$-stationary, we write $D(M\|L):=D_{KL}^{\pi}(M\|L)$ which can be interpreted as the KL divergence rate from $L$ to $M$, see Rached et al. (2004). Notably $f$-divergences of Markov chains have also been proposed for estimating the transition matrix from samples generated from Markov chains, for instance in Hao et al. (2018).
@@
This notion of product chain has appeared in (Levin and Peres, 2017, Exercise $12.7$) and differs from another slightly different “product-type chain” in Mathé (1998) or “product chain” introduced in (Levin and Peres, 2017, Section $12.4,20.4$). Analogously, a probability mass $\mu\in\mathcal{P}(\bigtimes_{i=1}^{d}\Omega^{(i)})$ is said to be of product form if there exists $\nu_{i}\in\mathcal{P}(\Omega^{(i)})$ for $i\in\llbracket d\rrbracket$ such that
@@
We remark that there exists a body of literature on various generalizations of mutual information to the case of $d>2$ random variables, see for example Ghassami and Kiyavash (2017) and the references therein.
@@
The Hellinger distance is commonly used to assess the convergence to equilibrium of product Markov chains, see for example Chen and Kumagai (2018); Levin and Peres (2017).
@@
Proof.
For brevity, throughout this proof we write $x=(x^{1},\ldots,x^{d})$ and $y=(y^{1},\ldots,y^{d})$. We first prove item (1). Since $D_{f}^{\pi}$ is a $f$-divergence from $\otimes_{i=1}^{d}L_{i}$ to $M$ with respect to $\pi$, it is non-negative according to Wang and Choi (2023). Since $\pi$ is a positive probability mass, equality holds if and only if for all $x\in\mathcal{X}$ we have $\widetilde{D}_{f}(M(x,\cdot);(\otimes_{i=1}^{d}L_{i})(x,\cdot))=0$ if and only if $M=\otimes_{i=1}^{d}L_{i}$ (see for instance Polyanskiy and Wu (2022)).
@@
where we apply (Levin and Peres, 2017, Lemma $20.9$) in the inequality. On the other hand, the lower bound can be seen via
@@
where the first inequality follows from (Chen and Kumagai, 2018, Proposition $2.3$). The desired result follows by taking maximum over $i\in\llbracket d\rrbracket$.
Finally, for item (5), we note that the proof is similar to (Choi and Wolfer, 2024, Section IIIA) and is therefore omitted. ∎

### v5

@@ Abstract
This paper analyzes the factorizability and geometry of transition matrices of multivariate Markov chains. Specifically, we demonstrate that the induced chains on factors of a product space can be regarded as information projections with respect to the Kullback-Leibler divergence. This perspective yields Han-Shearer type inequalities and submodularity of the entropy rate of Markov chains, as well as applications in the context of large deviations and mixing time comparison. As concrete algorithmic applications in Markov chain Monte Carlo (MCMC), we provide two illustrations based on lifted MCMC and swapping algorithm respectively to demonstrate projection samplers improve mixing over the original samplers. The projection sampler based on the swapping algorithm resamples the highest-temperature coordinate at stationarity at each step, and we prove that such practice accelerates the mixing time by multiplicative factors related to the number of temperatures and the dimension of the underlying state space when compared with the original swapping algorithm. Through simple numerical experiments on a bimodal target distribution, we show that the projection samplers mix effectively, in contrast to lifted MCMC and the swapping algorithm, which mix less well.
keywords:
Markov chains, $f$-divergence, mutual information, Markov chain Monte Carlo, large deviations, matrix nearness problem, swapping algorithm, Han’s inequality, Shearer’s lemma, submodularity, information geometry
{MSCcodes}
60F10, 60J10, 60J22, 94A15, 94A17
@@ 1 Introduction
(1) $\widetilde{D}_{KL}(\mu\|\nu):=\sum_{x\in\Omega}\mu(x)\ln\left(\dfrac{\mu(x)}{\nu(x)}\right),$
where the usual convention of $0\ln(0/0):=0$ applies. With the above setup in mind, the classical notion of mutual information between $X,Y$ is defined to be
@@
where the symbol $(\mu\otimes\nu)(x,y):=\mu(x)\nu(y)$ for all $x,y\in\Omega$ denotes the product distribution of $\mu$ and $\nu$. Note that $\mu\otimes\nu\in\mathcal{P}(\Omega^{2})$. It can be shown, via the chain rule of the KL divergence [9, Theorem $2.5.3$], that $p_{X}\otimes p_{Y}$ is the unique closest product distribution to $p_{(X,Y)}$ in the sense that
(2) $I(X;Y)=\min_{\mu,\nu\in\mathcal{P}(\Omega)}\widetilde{D}_{KL}(p_{(X,Y)}\|\mu\otimes\nu).$
In other words, the mutual information $I(X;Y)$ can broadly be interpreted as an entropic “distance” to the closest product distribution, that is, an entropic distance to independence. From (2), we immediately see that $I(X;Y)$ is non-negative, and vanishes if and only if $X,Y$ are independent [9, equation $(2.90)$].
In the context of Markov chains, we consider two transition matrices $M,L$ on a common finite state space $\mathcal{X}$ and we write the set of all transition matrices on $\mathcal{X}$ to be $\mathcal{L}(\mathcal{X})$. Let $f:\mathbb{R}^{+}\to\mathbb{R}$ be a convex function with $f(1)=0$, and $\pi\in\mathcal{P}(\mathcal{X})$. Analogous to $f$-divergence between infinitesimal generators of continuous-time Markov chains in [12, Proposition $1.5$], we define the $f$-divergence from $L$ to $M$ with respect to $\pi$ to be
@@
where $x^{i},y^{i}\in\mathcal{X}^{(i)}$ for $i=1,2$. Suppose now the state space $\mathcal{X}=\mathcal{X}^{(1)}\times\ldots\times\mathcal{X}^{(d)}$ takes on a product form for $d\in\mathbb{N}$. Given a $P\in\mathcal{L}(\mathcal{X})$, what is the closest product Markov chain? To put this question in concrete applications, we can think of an interacting particle system with $d$ particles or agents, such as the voter model [27], that is described by a transition matrix $P$. What is the dynamics of the closest independent system to $P$ in which each particle or agent evolves independently of each other? That is, we are interested in seeking a minimizer of
@@
which is analogous to the classical mutual information as in (2). Note that we write $\llbracket a,b\rrbracket:=\{a,a+1,\ldots,b\}$ for $a,b\in\mathbb{Z}$ and $\llbracket d\rrbracket:=\llbracket 1,d\rrbracket$ for $d\in\mathbb{N}$. We will first verify that $\mathbb{I}_{f}^{\pi}$ shares similar geometric properties with the distance to product distributions (i.e. mutual information between two random variables). As a result, $\mathbb{I}_{f}^{\pi}(P)$ can be interpreted as a distance to independence of a given $P$.
Orthogonality considerations allow us to identify and determine the closest product chain under KL divergence. This can be seen as a Markov chain version of the matrix nearness problem investigated in [26, 20], as we are seeking the closest product chain from a given $P$. In the dual case when $f$ generates the reverse KL divergence, we present a large deviation principle of Markov chains where $\mathbb{I}_{f}^{\pi}(P)$ plays a role in the exponent of large deviation probability. These results are presented in Section 2.1 and 2.1.1 below.
Note that in [24] a similar problem has been investigated in the context of diffusion processes, where the author studies the closest independent diffusion process of a given multivariate diffusion process and identifies the associated Wasserstein gradient flow and consequences for the McKean-Vlasov equation. On the other hand in the present manuscript, we focus on the closest independent Markov chain problem and the underlying geometry induced by information divergences between transition matrices.
We proceed to generalize these notions further. We introduce the leave-one-out, or more generally leave-$S$-out transition matrix, and investigate the factorizability of a transition matrix with respect to a partition or cliques of a given graph in Section 2.2 to 2.4. Observing that leave-$S$-out transition matrices are instances of Markov chain decomposition [23] or induced chains [1], we deduce comparison results for hitting and mixing time parameters such as spectral gap and log-Sobolev constant between $P$ and its information projections.
@@
We conclude this introduction by providing a simple motivating example in the context of lifted MCMC, where projection can yield improved sampler.
1.1 Motivating examples: lifted samplers
As a simple illustration to demonstrate the idea of projection samplers, we consider lifted MCMC samplers followed by projection to further improve mixing.
Precisely, consider a Metropolis-Hastings chain with transition matrix $Q=Q(M,\pi^{(1)})$ on the state space $\mathcal{X}^{(1)}$, where $M$ is the proposal chain and $\pi^{(1)}$ is the target distribution that we seek to sample from. For simplicity in this example we shall consider $\mathcal{X}^{(1)}=\llbracket-n,n\rrbracket$ for $n\in\mathbb{N}$.
To add memory and to avoid diffusive-like behaviour in the dynamics, one acceleration method is to consider the lifted Metropolis-Hastings chain with transition matrix $P$ on the augmented state space $\mathcal{X}=\mathcal{X}^{(1)}\times\{-1,+1\}$, where the second coordinate can now be interpreted as a direction or velocity variable. Specifically, we consider $P$ to be of run-and-tumble type [40]. From an initial state of $(x,v)\in\mathcal{X}$, $P$ moves according to the following rules:
  • (Position move) With probability $a$, $P$ moves from $(x,v)$ to $(y,v)$ according to $Q$.
  • With probability $b$, if $v=1$, $P$ moves from $(x,v)$ to $(\min\{x+1,n\},v)$. Similarly, if v=1v=-1, $P$ moves from $(x,v)$ to $(\max\{x-1,-n\},v)$.
  • (Flipping the direction) With probability $c$, $P$ moves from $(x,v)$ to (x,v)(x,-v).
We suppose that $a+b+c=1$. It can readily be shown that $P$ is $\pi=\pi^{(1)}\otimes\mathcal{U}(\{-1,+1\})$-stationary and is in general non-reversible with respect to $\pi$, where we denote $\mathcal{U}(\{-1,+1\})$ to be the discrete uniform distribution on the two-point space $\{-1,+1\}$. Such $P$ can be understood as a simplifed version of the kinetic random walks or Langevin diffusions [31], run-and-tumble models [40] or Gustafson’s guided walk samplers [17]. We are interested in comparing the following three MCMC samplers:
  • Original $Q$
  • Lifted sampler $P$, and we discard the samples associated with the direction coordinate
  • The keep-$\{1\}$-in projection sampler $P^{(1)}$, see Definition 2.18 below. To simulate one step of $P^{(1)}$, from an initial state of $x\in\mathcal{X}$, we draw uniformly at random a direction coordinate $v_{1}\sim\mathcal{U}(\{-1,+1\})$, then one moves from $(x,v_{1})$ to $(y,v_{2})$ according to $P$, followed by discarding $v_{2}$ and maintaining only $y$.
Comparing $Q$ and $P^{(1)}$, the latter can be understood as a suitably perturbed version of the former with moves that might have a small probability of taking place in $Q$. For instance, with probability $b$, $P^{(1)}$ is able to move from $x$ to $x+v$, in which such proposal move might have a much smaller probability in $Q$ than $b$.
Comparing $P$ and $P^{(1)}$, the intuitive rationale for the acceleration effect of $P$ lies in the added memory because of the inclusion of the velocity coordinate. For $P^{(1)}$, such memory seems to be lost as its velocity coordinate is randomized at each step according to $\mathcal{U}(\{-1,+1\})$.
The results established in this paper give that it is favourable to consider $P^{(1)}$ over $P$ for improved mixing: the KL divergence from $\Pi$ to $P$ is at least greater than or equal to that from $\Pi^{(1)}$ to $P^{(1)}$ ($\Pi$ or $\Pi^{(1)}$ are respectively the matrix where each row is $\pi$ or $\pi^{(1)}$), see Corollary 2.39. In addition, the multiplicative spectral gap of $P^{(1)}$ is at least as good as that of $P$, see Corollary 2.44.
1.1.1 Numerical experiments
For reproducibility, the code used in our experiments is available at https://github.com/mchchoi/factorization/tree/main. We first state the parameters of the experiments:
  • $\pi^{(1)}(x)\propto 2^{|x|}$. Such bimodal V-shaped target distribution is commonly used to assess the performance of MCMC samplers, see e.g. [11, 29]. Notably there are two modes at $\pm n$ respectively.
  • $n=20$.
  • $M$, the proposal chain, moves from $x$ to $\min\{x+1,n\}$ and $\max\{x-1,-n\}$ with probability $1/2$, and 0 otherwise.
  • $a=b=\frac{1}{4},c=\frac{1}{2}$.
  • All samplers are initialized at 20-20, the mode on the left, and are simulated for $1,000,000$ steps.
The results are summarized and presented in Figure 1, Table 1 and 2.
First, we note that $Q$ does not exhibit mixing: from the traceplot, histogram and empirical mean, it only explores the basin around the left mode at 20-20 and do not traverse to the right mode at $20$ in the experiment.
Second, from the traceplots and histograms we see that both $P$ and $P^{(1)}$ are able to hop between the two modes. One notable difference is that $P$ has more frequent hopping compared with $P^{(1)}$. While the histogram that these two generated are visually similar, from Table 1 the empirical distribution generated by $P^{(1)}$ is closer to the ground truth $\pi^{(1)}$ than that generated by $P$. From Table 2 the empirical mean and second moment generated by $P^{(1)}$ are also closer to the respective ground truth than that generated by $P$. These two tables seem to suggest that $P^{(1)}$ mixes better than the $x$-coordinate of $P$, thus offering empirical evidence that it is advantageous to use the projection sampler $P^{(1)}$ over either $P$ or $Q$ to sample from $\pi^{(1)}$.
(a) Traceplot and histogram of the trajectories of $Q$.
(b) Traceplot and histogram of the trajectories of the $x$-coordinate of $P$.
(c) Traceplot and histogram of the trajectories of $P^{(1)}$.
Figure 1: Numerical experiments comparing the three samplers $Q,P,P^{(1)}$ with target distribution being the V-shaped $\pi^{(1)}(x)\propto 2^{|x|}$.
Sampler $\widetilde{D}_{TV}(\widehat{\pi}^{(1)},\pi^{(1)})$ $\widetilde{D}_{KL}(\widehat{\pi}^{(1)}\|\pi^{(1)})$
Q 0.50 0.69
$P$ ($x$-only) 0.37 0.81
$P^{(1)}$ 0.31 0.53
Table 1: Comparison of total variation distance and KL divergence between $\widehat{\pi}^{(1)}$ and the ground truth $\pi^{(1)}$, where $\widehat{\pi}^{(1)}$ is the empirical distribution formed by the trajectories of the samplers. Recall that $\widetilde{D}_{TV}(\mu,\nu):=\tfrac{1}{2}\sum_{x}|\mu(x)-\nu(x)|$ and $\widetilde{D}_{KL}$ is defined in (1).
Sampler Mean Second moment
Q -19.00 362.93
$P$ ($x$-only) -4.44 284.81
$P^{(1)}$ -2.60 301.07
Truth $\pi^{(1)}$ 0 363.00
Table 2: Comparison of the first and second moment between the samplers and the ground truth $\pi^{(1)}$.
1.2 Organization of the paper
@@
The rest of this paper is organized as follows. In Section 2, we first recall the notion of $f$-divergences between transition matrices and probability measures. We derive a few important properties of these divergences on finite product state spaces, which allow us to define an entropic distance to independence of a given multivariate $P$ in Section 2.1. In Section 2.1.1, we determine and identify the closest product chain under KL divergence, and present a large deviation principle in this context. We investigate the factorizability of $P$ with respect to partition or cliques of a given graph in Section 2.2 to 2.4. In Section 2.5, we compare mixing and hitting time parameters between $P$ and its information projections, while in Section 2.6, we show that several entropic functions that naturally arise in this paper are in fact submodular. To illustrate the applicability of projection chains, in Section 3 we propose a projection sampler and compare its mixing time with the original swapping algorithm, along with some simple numerical experiments in Section 3.3.
@@ 2 Distance to independence and factorizability of Markov chains
On a finite state space $\mathcal{X}$, we define $\mathcal{L}=\mathcal{L}(\mathcal{X})$ as the set of transition matrices of discrete-time homogeneous Markov chains. We denote by $\mathcal{P}(\mathcal{X})$ to be the set of probability masses on $\mathcal{X}$. Let $\pi\in\mathcal{P}(\mathcal{X})$ be any given positive probability distribution (i.e. $\pi$ satisfies $\min_{x}\pi(x)>0$), and denote $\mathcal{L}(\pi)\subset\mathcal{L}$ as the set of $\pi$-reversible transition matrices on $\mathcal{X}$, where a transition matrix $P\in\mathcal{L}$ is said to be $\pi$-reversible if $\pi(x)P(x,y)=\pi(y)P(y,x)$ for all $x,y\in\mathcal{X}$. We also say that $P\in\mathcal{L}$ is $\pi$-stationary if it satisfies $\pi P=\pi$. Suppose that $P$ is $\pi$-stationary, then the $\pi$-dual or the time reversal of $P$, $P^{*}$, is defined to be $P^{*}(x,y):=\frac{\pi(y)}{\pi(x)}P(y,x)$, for all $x,y\in\mathcal{X}$.
@@
(3) $D_{f}^{\pi}(M\|L):=\sum_{x\in\mathcal{X}}\pi(x)\sum_{y\in\mathcal{X}}L(x,y)f\left(\dfrac{M(x,y)}{L(x,y)}\right).$
@@
(4) $\widetilde{D}_{f}(\mu\|\nu):=\sum_{x\in\Omega}\nu(x)f\left(\frac{\mu(x)}{\nu(x)}\right),$
where we apply the usual convention that $0f(\frac{0}{0}):=0$ and $0f(\frac{a}{0}):=af^{\prime}(+\infty)$ with $f^{\prime}(+\infty):=\lim_{x\to 0^{+}}xf(\frac{1}{x})$ for $a>0$ in the two definitions above. We also adapt the convention that $0\cdot\infty:=0$.
In the special case of taking $f(t)=t\ln t$, we recover the KL divergence. In this case we shall write $D_{KL}^{\pi}$ and $\widetilde{D}_{KL}$ respectively. In particular, when $M,L$ are assumed to be $\pi$-stationary, we write $D(M\|L):=D_{KL}^{\pi}(M\|L)$ which can be interpreted as the KL divergence rate from $L$ to $M$, see [37]. Notably $f$-divergences of Markov chains have also been proposed for estimating the transition matrix from samples generated from Markov chains, for instance in [18].
@@
This notion of product chain has appeared in [25, Exercise $12.7$] and differs from another slightly different “product-type chain” in [30] or “product chain” introduced in [25, Section $12.4,20.4$]. Analogously, a probability mass $\mu\in\mathcal{P}(\bigtimes_{i=1}^{d}\Omega^{(i)})$ is said to be of product form if there exists $\nu_{i}\in\mathcal{P}(\Omega^{(i)})$ for $i\in\llbracket d\rrbracket$ such that
@@
We remark that there exists a body of literature on various generalizations of mutual information to the case of $d>2$ random variables, see for example [16] and the references therein.
@@
The Hellinger distance is commonly used to assess the convergence to equilibrium of product Markov chains, see for example [6, 25].
@@
Proof 2.6.
For brevity, throughout this proof we write $x=(x^{1},\ldots,x^{d})$ and $y=(y^{1},\ldots,y^{d})$. We first prove item (1). Since $D_{f}^{\pi}$ is a $f$-divergence from $\otimes_{i=1}^{d}L_{i}$ to $M$ with respect to $\pi$, it is non-negative according to [43]. Since $\pi$ is a positive probability mass, equality holds if and only if for all $x\in\mathcal{X}$ we have $\widetilde{D}_{f}(M(x,\cdot);(\otimes_{i=1}^{d}L_{i})(x,\cdot))=0$ if and only if $M=\otimes_{i=1}^{d}L_{i}$ (see for instance [36]).
@@
where we apply [25, Lemma $20.9$] in the inequality. On the other hand, the lower bound can be seen via
@@
where the first inequality follows from [6, Proposition $2.3$]. The desired result follows by taking maximum over $i\in\llbracket d\rrbracket$.
Finally, for item (5), we note that the proof is similar to [8, Section IIIA] and is therefore omitted.